Nova Company’s total overhead cost at various levels of activity are presented below:

Month Machine-
Hours
Total
Overhead
Cost
April 48,000 $ 193,480
May 38,000 $ 169,880
June 58,000 $ 217,080
July 68,000 $ 240,680

Assume that the total overhead cost above consists of utilities, supervisory salaries, and maintenance. The breakdown of these costs at the 38,000 machine-hour level of activity is:

Utilities (variable) $ 45,600
Supervisory salaries (fixed) 67,000
Maintenance (mixed) 57,280
Total overhead cost $ 169,880

Nova Company’s management wants to break down the maintenance cost into its variable and fixed cost elements.

Required:

1. Estimate how much of the $240,680 of overhead cost in July was maintenance cost. (Hint: to do this, it may be helpful to first determine how much of the $240,680 consisted of utilities and supervisory salaries. Think about the behavior of variable and fixed costs.)

2. Using the high-low method, estimate a cost formula for maintenance in the form Y = a + bX.

3. Express the company’s total overhead cost in the form Y = a + bX.

4. What total overhead cost would you expect to be incurred at an activity level of 43,000 machine-hours?
Answer #1

1) Calculate July Maintenance cost

Total July Cost 240680
Utilities (45600/38000*68000) 81600
Supervisory salaries 67000
Maintenance cost (240680-81600-67000) 92080

2) variable cost per unit = (92080-57280)/(68000-38000) = 1.16 per unit

Fixed cost = 57280-(38000*1.16) = 13200

Y = a+bx

Y = 13200+1.16X

3) Y = a+bx

Y = 80200+2.36X

4) Total overhead cost = 80200+(43000*2.36) = 181680